# LLM-Ensemble: Optimal Large Language Model Ensemble Method for E-commerce Product Attribute Value Extraction

### v1

@@ Abstract.
In this paper, we propose a novel algorithm called LLM-ensemble to ensemble different LLMs’ outputs for attribute value extraction. We iteratively learn the weights for different LLMs to aggregate the labels with weights to predict the final attribute value. Not only can our proposed method be proven theoretically optimal, but it also ensures efficient computation, fast convergence, and safe deployment. We have also conducted extensive experiments with various state-of-the-art LLMs, including Llama2-13B, Llama2-70B, PaLM-2, GPT-3.5, and GPT-4, on Walmart’s internal data. Our offline metrics demonstrate that the LLM-ensemble method outperforms all the state-of-the-art single LLMs on Walmart’s internal dataset. This method has been launched in several production models, leading to improved Gross Merchandise Volume (GMV), Click-Through Rate (CTR), Conversion Rate (CVR), and Add-to-Cart Rate (ATC).
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: The 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington D.C., USAisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing E-commerce infrastructureccs: Computing methodologies Machine learning
@@ 2.1. Product Attribute Value Extraction
Early research on attribute value extraction relied on domain-specific rules to identify attribute-value pairs from product descriptions, as indicated in (Zhang et al., 2009; Vandic et al., 2012). However, the initial learning-based approaches required substantial feature engineering and struggled to adapt to previously unseen attributes and values (Ghani et al., 2006; Putthividhya and Hu, 2011; Wong et al., 2009). With the advancement of neural networks, more recent studies (Kozareva et al., 2016; Zheng et al., 2018) have shifted towards employing BiLSTM-CRF architectures for tagging attribute values in product titles. OpenTag (Zheng et al., 2018) leverages a BiLSTM-CRF model enhanced by active learning. Further innovations include SU-OpenTag (Xu et al., 2019), which extends OpenTag by incorporating both a target attribute and the product title into a pre-trained language model. AdaTag (Yan et al., 2021) introduces a combination of a language model and a mixture-of-experts module for attribute value extraction, while TXtract (Karamanolakis et al., 2020) integrates a product taxonomy into its model. Approaches like AVEQA (Wang et al., 2020) and MAVEQA (Yang et al., 2022) frame attribute value extraction as a question-answering task, utilizing diverse pre-trained language models to process the target attribute, product category, and title. OA-Mine (Zhang et al., 2022) explores the mining of unknown attribute values and attributes using a language model. More recent research has explored soft prompt tuning to fine-tune a minimal number of trainable parameters within a language model (Blume et al., 2023; Yang et al., 2023). Additionally, Brinkmann et al. (Brinkmann et al., 2023) have demonstrated the extraction of product attribute values using Large Language Models (LLMs) like GPT-3.5 and GPT-4, highlighting the ongoing evolution in this field.
@@ 3.2. LLM Ensemble
Initialization: $v_{qi}=1,\forall i\in\mathcal{N};T_{qij}=I(W_{qij}\neq 0),\forall i\in\mathcal{N},\forall j\in\mathcal{P}$  // I is the indicator matrix
1
@@
3       $\hat{y}_{qj}\leftarrow\operatorname*{argmax}\sum_{i=1}^{N}v_{qi}I(W_{qij}=k),\forall j\in\mathcal{P}$ $\hat{\alpha}_{qi}\leftarrow\frac{\sum_{j=1}^{P}I(W_{qij}=\hat{y}_{qj})}{\sum_{j=1}^{P}T_{qij}},\forall i\in\mathcal{N}$ $v_{qi}\leftarrow L\hat{\alpha}_{qi}-1,\forall i\in\mathcal{N}$
@@ 4.1. Comparison Experiments
Our proposed LLM-ensemble is compared with its base LLMs: Llama2-13B, Llama2-70B (Touvron et al., 2023), PaLM-2 (Chowdhery et al., 2023), GPT-3.5, GPT-4 (Brown et al., 2020). We also include logistic regression (Hosmer Jr et al., 2013) and our internal rule-based method in the baseline models. The datasets we employ are Walmart-Age and Walmart-Gender, which contain products sensitive to the ages or genders of customers [^1] Due to the privacy policy in Walmart, we only disclose these two attributes. The actual product in practice has more attributes involved.. Each dataset has 20M items for offline evaluation. The ground-truth label is created by the crowdsourcing results of human experts.
@@ 5. Conclusion
this paper introduces an innovative ensemble method, LLM-ensemble, designed to optimize product attribute value extraction by ensembling the outputs of various Large Language Models (LLMs). By dynamically learning the weights for different LLMs and aggregating their labels, our algorithm not only achieves theoretical optimality but also excels in efficiency, convergence speed, and deployment safety. Through comparison experiments with state-of-the-art LLMs on Walmart’s internal dataset, the LLM-ensemble method has demonstrated superior performance over all individual LLMs. The practical application of this method in production models has notably enhanced Gross Merchandise Volume (GMV), Click-Through Rate (CTR), Conversion Rate (CVR), and Add-To-Cart Rate (ATC), thereby contributing significantly to the advancement of NLP applications in e-commerce and setting a new benchmark for high-quality recommendations and customer satisfaction.

### v2

@@ Abstract.
In this paper, we propose a novel algorithm called LLM-ensemble to ensemble different LLMs’ outputs for attribute value extraction. We iteratively learn the weights for different LLMs to aggregate the labels with weights to predict the final attribute value. Not only can our proposed method be proven theoretically optimal, but it also ensures efficient computation, fast convergence, and safe deployment. We have also conducted extensive experiments with various state-of-the-art LLMs on Walmart’s internal data. Our offline metrics demonstrate that the LLM-ensemble method outperforms all the state-of-the-art single LLMs on Walmart’s internal dataset. This method has been launched in several production models, leading to improved Gross Merchandise Volume (GMV), Click-Through Rate (CTR), Conversion Rate (CVR), and Add-to-Cart Rate (ATC).
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/3626772.3661357ccs: Applied computing E-commerce infrastructureccs: Computing methodologies Machine learning
@@ 2.1. Product Attribute Value Extraction
Early research on attribute value extraction relied on domain-specific rules to identify attribute-value pairs from product descriptions, as indicated in (Zhang et al., 2009; Vandic et al., 2012). However, the initial learning-based approaches required substantial feature engineering and struggled to adapt to previously unseen attributes and values (Ghani et al., 2006; Putthividhya and Hu, 2011; Wong et al., 2009). Recently, some studies (Kozareva et al., 2016; Zheng et al., 2018) have shifted towards employing BiLSTM-CRF architectures for tagging attribute values in product titles. OpenTag (Zheng et al., 2018) leverages a BiLSTM-CRF model enhanced by active learning. Further innovations include SU-OpenTag (Xu et al., 2019), which extends OpenTag by incorporating both a target attribute and the product title into a pre-trained language model. AdaTag (Yan et al., 2021) introduces a combination of a language model and a mixture-of-experts module for attribute value extraction, while TXtract (Karamanolakis et al., 2020) integrates a product taxonomy into its model. Approaches like AVEQA (Wang et al., 2020) and MAVEQA (Yang et al., 2022) frame attribute value extraction as a question-answering task, utilizing diverse pre-trained language models to process the target attribute, product category, and title. OA-Mine (Zhang et al., 2022) explores the mining of unknown attribute values and attributes using a language model. More recent research has explored soft prompt tuning to fine-tune a minimal number of trainable parameters within a language model (Blume et al., 2023; Yang et al., 2023). Additionally, Brinkmann et al. (Brinkmann et al., 2023) have demonstrated the extraction of product attribute values using Large Language Models (LLMs) like GPT-3.5 and GPT-4, highlighting the ongoing evolution in this field. Zou et al. (Zou et al., 2024a, b) propose ImplicitAVE and EIVEN to extract implicit attribute values with multimodal large language models.
@@ 3.2. LLM Ensemble
Initialization: $v_{qi}=1,\forall i\in\mathcal{N};T_{qij}=I(W_{qij}\neq 0),\forall i\in\mathcal{N},\forall j\in\mathcal{P}$
  // I is the indicator matrix
2
@@
4       $\hat{y}_{qj}\leftarrow\operatorname*{argmax}\sum_{i=1}^{N}v_{qi}I(W_{qij}=k),\forall j\in\mathcal{P}$
5       $\hat{\alpha}_{qi}\leftarrow\frac{\sum_{j=1}^{P}I(W_{qij}=\hat{y}_{qj})}{\sum_{j=1}^{P}T_{qij}},\forall i\in\mathcal{N}$
6       $v_{qi}\leftarrow L\hat{\alpha}_{qi}-1,\forall i\in\mathcal{N}$
@@ 4.1. Comparison Experiments
Our proposed LLM-ensemble is compared with its base LLMs: Llama2-13B, Llama2-70B (Touvron et al., 2023), PaLM-2 (Chowdhery et al., 2023), GPT-3.5, GPT-4 (Brown et al., 2020). We also include logistic regression (Hosmer Jr et al., 2013) and our internal rule-based method in the baseline models. The datasets we employ are Walmart-Age and Walmart-Gender, which contain products sensitive to the ages or genders of customers [^1] Due to the privacy policy in Walmart, we only disclose these two attributes. The actual product in practice has more attributes involved.. Each dataset has 20K items for offline evaluation. The ground-truth label is created by the crowdsourcing results.
@@ 5. Conclusion
this paper introduces an innovative ensemble method, LLM-ensemble, designed to optimize product attribute value extraction by ensembling the outputs of various Large Language Models (LLMs). By dynamically learning the weights for different LLMs and aggregating their labels, our algorithm not only achieves theoretical optimality but also excels in efficiency, convergence speed, and deployment safety. Through comparison experiments with state-of-the-art LLMs on Walmart’s internal dataset, the LLM-ensemble method has demonstrated superior performance over all individual LLMs. The A/B test of this method in production models has also improved the performance of our recommendation product.